1. (3.) Using the definition of an inverse, are f(x) and g(x) inverses of one another? Why or why not?
1
Expert's answer
2021-09-13T07:54:26-0400

Let consider these function

Let


"f(x)=\frac{x+3}{2}"


"g(x)=2x-3"

If function "f(x)" and "g(x)" are inverses, their composition will equal "x"


Composition 1:

"f(g(x))=f(x)=\frac{(2x+3)+3}{2}=\frac{2x}{2}=x"


Composition 2:

"g(f(x))
=2(\frac{x+3}{2})-3=x+3-3=x"


Hence these are inverse as their compositions equal x.